# ON THE FACES OF UNIGRAPHIC $3$-POLYTOPES

### v2

@@ Abstract
R.W. Maffucci, University of Coventry, United Kingdom CV1
@@ 2.2   Setup
$\exists k:1\leq k\leq n,\quad F_{k+1}=[u_{k+1},u_{k+2},x_{k}],\qquad\deg(u_{k+1})=\deg(u_{k+2})=3,$
@@
both external to the face $F$. By planarity, the paths (2.13) must intersect at a vertex $v_{1}$. Since the faces $F_{k},F_{3},F_{\xi(k)},F_{\xi(3)}$ are pairwise disjoint, it follows that $P_{5}$ and $P_{7}$ intersect at $v_{1}$. Likewise, $P_{5},P_{8}$ intersect at a vertex $v_{2}$, $P_{6},P_{7}$ intersect at a vertex $v_{3}$, and $P_{6},P_{8}$ intersect at a vertex $v_{4}$. Recall that $P_{5},P_{6}$ are internally disjoint, and $P_{7},P_{8}$ are internally disjoint. Therefore, $F^{\prime}_{k}\cap F^{\prime}_{3}$ contains at least four vertices, thus in fact $F^{\prime}_{k}=F^{\prime}_{3}=:F^{\prime}$, so that there exists a face $F^{\prime}\neq F$ of $G$ containing all of $x_{k},x_{3},y_{k},y_{3}$.
@@
With the above context and notation, assume that $u_{a},u_{i},u_{b},u_{j},u_{c},u_{k}$ appear around boundary of $F$ in this order, where $1\leq i<j<k\leq n$,
@@
We claim that at least three pairwise non-consecutive elements of (2.15) contain $z$. We argue by contradiction. Since three consecutive elements of (2.15) cannot all contain $z$, and moreover there do not exist three pairwise non-consecutive elements of (2.15) containing $z$, then in fact at most four elements of (2.15) may contain $z$, and given any three of them, exactly two of them are consecutive. W.l.o.g., $z\not\in F_{3},F_{4},F_{n}$. Similarly to Lemma 2.5, we have $\mathcal{T}_{1}(F,n,3)(G)\simeq\mathcal{T}_{1}(F,n,4)(G)\simeq G$, leading to either $n-3=n-4$, or $n-3=4$, contradiction.
@@
$|i-j|,\ |j-k|,\ |i-k|$
equals $2$ or less, thus it is exactly $2$. We may then write w.l.o.g.
$z\in F_{1},F_{3},F_{a-1},\quad z\not\in F_{2},F_{4},F_{a},\qquad 6\leq a\leq n,$
where $a$ is the smallest index $>4$ such that $z\not\in F_{a}$. On one hand, at least one of $F_{5},F_{6},F_{7}$ does not contain $z$. On the other hand, assuming $z\not\in F_{7}$, similarly to Lemma 2.5, we obtain $\mathcal{T}_{1}(F,7,2)(G)\simeq\mathcal{T}_{1}(F,7,4)(G)\simeq G$, hence either $7-2=7-4$ or $7-2=n-(7-4)$, contradiction. Hence $a=6$ and $z\in F_{7}$.
Likewise, at least one of $F_{7},F_{8},F_{9}$ does not contain $z$. Assuming $z\not\in F_{9}$ leads to $\mathcal{T}_{1}(F,9,4)(G)\simeq\mathcal{T}_{1}(F,9,6)(G)\simeq G$, i.e. again either $5=3$ or $n=8$, contradiction. Hence $z\not\in F_{8}$ and $z\in F_{9}$. It now follows that $\mathcal{T}_{1}(F,8,4)(G)\simeq\mathcal{T}_{1}(F,8,2)(G)\simeq G$, whence $n=10$. The three consecutive $F_{9},F_{10},F_{1}$ cannot all contain $z$, thus $z\not\in F_{10}$. Summarising,
@@
Looking at the isomorphism $\mathcal{T}_{1}(F,2,6)$, we conclude that $F_{2},F_{6}$ are a $4$-gon and a $6$-gon in some order. If $F_{2}$ is a $4$-gon and $F_{6}$ a $6$-gon (the other case is analogous), we analyse in turn the isomorphisms $\mathcal{T}_{1}(F,6,10)$, $\mathcal{T}_{1}(F,10,4)$, $\mathcal{T}_{1}(F,4,8)$, and $\mathcal{T}_{1}(F,8,2)$, to see that $F_{10}$ is a $4$-gon, $F_{4}$ a $6$-gon, $F_{8}$ a $4$-gon, and $F_{2}$ a $6$-gon, contradiction. ∎
@@
Now assume instead that (2.2) holds, so that fixing w.l.o.g. $k=1$ and defining $x:=x_{1}$ we have
$F_{2}=[u_{2},u_{3},x],\qquad\deg(u_{2})=\deg(u_{3})=3,$
@@
We repeat the above steps, where the roles of $u_{2},u_{3}$ are swapped, the roles of $u_{1},u_{4}$ are swapped, and the roles of $u_{n},u_{5}$ are swapped. Hence $G[5,4,3]$ is a polyhedron of same sequence as $G$. If it is not isomorphic to $G$, the proof is complete. If it is, we remark that $G[5,4,3]\not\simeq G[5,4,2]$, hence $G[5,4,2]$ is a simple, planar graph not isomorphic to $G$, and of same sequence $\sigma$. If $G[5,4,2]$ is $3$-connected, the proof is complete. Otherwise, in $G$ the faces $F_{1}$ and $F_{4}$ have a common vertex $z^{\prime}$, where $z^{\prime}\neq x$. Note that $F_{3}$ and $F_{n}$ have a common vertex $z$, and $F_{1}$ and $F_{4}$ have a common vertex $z^{\prime}$, hence by planarity of $G$ we deduce that $z=z^{\prime}$.
@@
However, the vertex $z$ cannot have degree $3$ in $G$, since the faces $F_{4},F_{n}$ are distinct. Therefore, one of $G[n,1,2],G[n,1,3],G[5,4,3],G[5,4,2]$ is a polyhedron of same degree sequence as $G$. The proof of Theorem 1.1 is complete.

### v3

@@ Abstract
R.W. Maffucci, Dipartimento di Matematica, Università di Torino
Via Carlo Alberto 10, Turin 10123, Italy
@@ 2.2   Setup
$\exists x\in V(G),\ \exists k:1\leq k\leq n,\quad F_{k+1}=[u_{k+1},u_{k+2},x],\qquad\deg(u_{k+1})=\deg(u_{k+2})=3,$
@@
both external to the face $F$. By planarity, the paths (2.13) must intersect at a vertex $v_{1}$. Since the faces $F_{k},F_{3},F_{\xi(k)},F_{\xi(3)}$ are pairwise disjoint, it follows that $P_{5}$ and $P_{7}$ intersect at $v_{1}$. Likewise, $P_{5},P_{8}$ intersect at a vertex $v_{2}$, $P_{6},P_{7}$ intersect at a vertex $v_{3}$, and $P_{6},P_{8}$ intersect at a vertex $v_{4}$. Recall that $P_{5},P_{6}$ are internally disjoint, and $P_{7},P_{8}$ are internally disjoint. Therefore, $F^{\prime}_{k}\cap F^{\prime}_{3}$ contains four distinct vertices $v_{1},v_{2},v_{3},v_{4}$, however two distinct faces may intersect in at most two vertices, thus in fact $F^{\prime}_{k}=F^{\prime}_{3}=:F^{\prime}$, so that there exists a face $F^{\prime}\neq F$ of $G$ containing all of $x_{k},x_{3},y_{k},y_{3}$.
@@
With the above context and notation, assume that $u_{a},u_{i},u_{b},u_{j},u_{c},u_{k}$ appear around the boundary of $F$ in this order, where $1\leq i<j<k\leq n$,
@@
We claim that at least three pairwise non-consecutive elements of (2.15) contain $z$. We argue by contradiction. Since three consecutive elements of (2.15) cannot all contain $z$, and moreover there do not exist three pairwise non-consecutive elements of (2.15) containing $z$, then in fact at most four elements of (2.15) may contain $z$, and given any three of them, exactly two of them are consecutive. W.l.o.g., $z\not\in F_{3},F_{4},F_{n}$, and $z\in F_{2},F_{b}$ with $15\leq b\leq n-$ 1. By the first part of the proof of Lemma 2.5, $\mathcal{T}_{1}(F,3,n)(G)\simeq G$, so that one of $F_{3}$, $F_{n}$ is a triangle and the other an $n-3$-gon, and similarly $\mathcal{T}_{1}(F,4,n)(G)\simeq G$, so that one of $F_{4}$, $F_{n}$ is a quadrangle and the other an $n-4$-gon. This leads to either $n-3=n-4$ or $n-3=4$, contradiction.
@@
$|i-j|,\ |j-k|,\ |i-k|$ (2.16)
equals $2$ or less, thus it is exactly $2$.
We consider two cases: either exactly two of (2.16) equal $2$, or exactly one of them equals $2$. In the first case, we may then write w.l.o.g.
$z\in F_{n-1},F_{1},F_{3},\quad z\not\in F_{n},F_{2},F_{4}.$
Assuming $z\not\in F_{5}$, we apply Lemma 2.5 with $a=1$, $i=2$, $b=3$, $j=5$, $c=n-1$, and $k=n$, to obtain $k-j=j-i$ thus $n=8$, contradiction. Hence $z\in F_{5}$.
Likewise, assuming $z\not\in F_{7}$, we apply Lemma 2.5 with $a=1$, $i=2$, $b=5$, $j=7$, $c=n-1$, and $k=n$, to obtain $n-7=7-2$. We also apply it with $a=3$, $i=4$, $b=5$, $j=7$, $c=n-1$, and $k=n$, to obtain $n-7=7-4$, contradicting $n-7=7-2$. Hence $z\in F_{7}$.
At least one of the three consecutive $F_{5},F_{6},F_{7}$ does not contain $z$, hence $z\not\in F_{6}$. We apply Lemma 2.5 with $a=1$, $i=2$, $b=5$, $j=6$, $c=n-1$, and $k=n$, obtaining $n=10$. The three consecutive $F_{7},F_{8},F_{9}$ cannot all contain $z$, thus $z\not\in F_{8}$. Summarising,
@@
Looking at the isomorphism $\mathcal{T}_{1}(F,2,6)$, we conclude that $F_{2},F_{6}$ are a $4$-gon and a $6$-gon in some order. If $F_{2}$ is a $4$-gon and $F_{6}$ a $6$-gon (the other case is analogous), we analyse in turn the isomorphisms $\mathcal{T}_{1}(F,6,10)$, $\mathcal{T}_{1}(F,10,4)$, $\mathcal{T}_{1}(F,4,8)$, and $\mathcal{T}_{1}(F,8,2)$, to see that $F_{10}$ is a $4$-gon, $F_{4}$ a $6$-gon, $F_{8}$ a $4$-gon, and $F_{2}$ a $6$-gon, contradiction.
It remains to inspect the second case, where exactly one of (2.16) equals $2$. We write w.l.o.g.
$z\in F_{n-1},F_{1},F_{m-1},\quad z\not\in F_{n},F_{2},F_{m},$
where $35\leq m\leq n-$ 3 otherwise we would be in the already analysed first case. Applying Lemma 2.5 with $a=1$, $i=2$, $b=m-1$, $j=m$, $c=n-1$, and $k=n$, we get $n-m=m-2$ thus $m=n/2+1$, and in particular $m\geq 6$.
Assuming $z\not\in F_{3}$, we apply Lemma 2.5 with $a=1$, $i=3$, $b=n/2$, $j=n/2+1$, $c=n-1$, and $k=n$, to obtain $n-(n/2+1)=(n/2+1)-3$, contradiction. Hence $z\in F_{3}$.
Next, assuming $z\not\in F_{5}$, we apply Lemma 2.5 with $a=1$, $i=2$, $b=3$, $j=5$, $c=n-1$, and $k=n$, to obtain $n=8$, contradiction. Hence $z\in F_{5}$.
At least one of the three consecutive $F_{3},F_{4},F_{5}$ does not contain $z$, hence $z\not\in F_{4}$. Now if $n=10$ we conclude as in the previous case. Otherwise, we apply Lemma 2.5 with $a=3$, $i=4$, $b=n/2$, $j=n/2+1$, $c=n-1$, and $k=n$, to obtain $n-(n/2+1)=(n/2+1)-4$, contradiction. The proof of Proposition 2.2 is complete. ∎
@@
Now assume instead that (2.2) holds, so that fixing w.l.o.g. $k=1$ we have
$\exists x\in V(G):\quad F_{2}=[u_{2},u_{3},x],\qquad\deg(u_{2})=\deg(u_{3})=3,$
@@
We apply the above discussion of the proof of $G[n,1,2]$ being a polyhedron where the roles of $u_{2},u_{3}$ are swapped, the roles of $u_{1},u_{4}$ are swapped, and the roles of $u_{n},u_{5}$ are swapped. Hence $G[5,4,3]$ is a polyhedron of same sequence as $G$. If it is not isomorphic to $G$, the proof is complete. If it is, we remark that $G[5,4,3]\not\simeq G[5,4,2]$, hence $G[5,4,2]$ is a simple, planar graph not isomorphic to $G$, and of same sequence $\sigma$.
We will now show that $G[5,4,2]$ is $3$-connected, completing the proof. Assume by contradiction that $G[5,4,2]$ is not 3-connected. We apply the above discussion, where we showed that if $G[n,1,3]$ is not 3-connected, then $F_{3}$ and $F_{n}$ have a common vertex $z\neq x$, but we swap the roles of $u_{2},u_{3}$, of $u_{1},u_{4}$, and of $u_{n},u_{5}$. Hence in $G$ the faces $F_{1}$ and $F_{4}$ have a common vertex $z^{\prime}$, where $z^{\prime}\neq x$. Note that $F_{3}$ and $F_{n}$ have a common vertex $z$, and $F_{1}$ and $F_{4}$ have a common vertex $z^{\prime}$, hence by planarity of $G$ we deduce that $z=z^{\prime}$.
@@
However, the vertex $z$ cannot have degree $3$ in $G$, since the faces $F_{4},F_{n}$ are distinct, contradiction. Hence, $G[5,4,2]$ is 3-connected, and thus a polyhedron.
Therefore, one of $G[n,1,2],G[n,1,3],G[5,4,3],G[5,4,2]$ is a polyhedron of same degree sequence as $G$. The proof of Theorem 1.1 is complete.
@@ Acknowledgements.
R. M. was partially supported by Programme for Young Researchers ‘Rita Levi Montalcini’ Discrete and Probabilistic Methods in Mathematics with Applications, awarded to R. M.